OpenStudy (anonymous):

Find the equation of the line that goes through the points (-2, -1) and (2, -5).

OpenStudy (jim766):

to write the equation of a line you need a slope and a y intercept, Do you know how to find the slope?

OpenStudy (anonymous):

y=mx+b

OpenStudy (jim766):

thats the slope intercept form of a line. Slope formula given two points is y2-y1/x2-x1 -5 --1/2 --1 tell me what you get

OpenStudy (anonymous):

3/2

OpenStudy (jim766):

-5 - -1 -5 +1 = -4 minus a neg, is same as addition ________ 2 - - 2 2 + 2 = 4 slope is -4/4, = -1 do you see where I got that?

OpenStudy (anonymous):

Yes, I didn't see the - in front of the 5

OpenStudy (jim766):

now that you have a slope, how do you find the equation? using point slope or y = mx + b? your choice

OpenStudy (anonymous):

y=mx+b

OpenStudy (jim766):

pick one of the points doesnt matter which one....which ever..

OpenStudy (jim766):

x y say (-2, -1), subsitute this and slope into y = mx + b -1 = -1(-2) + b , b is the y intercept, solve for b

OpenStudy (anonymous):

(-2,-1) y=mx+b -1=-1(-2)+b -1=2+b -2 -2 -3=b b=-3

OpenStudy (jim766):

great! now you have a slope and a y intercpet for your line. m = -1, b = -3, just substitute them into y = mx + b y = -1x - 3 or y = -x -3

OpenStudy (anonymous):

ok. Thanks